OpenStudy (anonymous):

Trig...What would the equation of the function be if the graph of the function y = 2 sin (4x) is shifted left 45 degrees and down 2?

OpenStudy (dumbcow):

45 degrees = pi/4 by shifting left, this means the "starting" point is now x=-pi/4 4(-pi/4) +C = 0 --> C = pi \[y = 2\sin(4x+\pi) -2\]

OpenStudy (dumbcow):

you can also show the shift this way: 4(x+pi/4) = 4x +pi